hello i need help with this one please
two cars start from the same place traveling in opposite directions.one car travels 4 miles per hour faster than the other car.find the speeds if after 5 hours they are 530 miles apart. thanks

Answer by stanbon(75887)

i need help with this one please
Two cars start from the same place traveling in opposite directions.
One car travels 4 miles per hour faster than the other car.
Find the speeds if after 5 hours they are 530 miles apart.
--------------
1st Car DATA:
time = 5 hrs ; rate = x mph ; distance = rt = 5x miles
----
2nd Car DATA:
time = 5 hrs ; rate = x+4 mph ; distance = rt = 5(x+4) miles
======================
Equation:
miles + miles = 530 miles
5x + 5(x+4) = 530
5x + 5x + 20 = 530
10x = 510
x = 51 mph (one of the cars)
x+4 = 55 mph (the other car)
================================
Cheers,
Stan H.
======================